the square of the hypotenuse of this triangle is 400 cm. Determine the actual measures of the hypotenuse and sides a and b of this right triangle.
Found 3 solutions by rfer, ewatrrr, Alan3354:
Answer by rfer(16322)

12, 16, 20 triangle

Answer by ewatrrr(24785)


Hi,

square of the hypotenuse of this triangle is 400 cm
Determine the actual measures of the hypotenuse and sides a and b of this right triangle
using the Pythagorean theorem
(20m)^2 = a^2 + b^2
400 m^2 = a^2 + b^2
IF a = 16m and b = 12m
THEN: 400 m^2 = 256 m^2 + 144 m^2

Answer by Alan3354(69443)

the square of the hypotenuse of this triangle is 400 cm. Determine the actual measures of the hypotenuse and sides a and b of this right triangle
----------------
Insufficient info.
It could be 20, 16 & 12, or
20, 10 and sqrt(300), or any of an ifinite # of combinations.